A beam of light is propagating in vacuum and its frequency is constant. Show that 

the average energy carried by it per unit area is proportional to the associated 

electric field vector.


1
Expert's answer
2021-06-03T09:16:07-0400

It can be shown energy per unit volume associated with electric field is "\frac{\varepsilon E^2}{2}" when the field is oscillating the average becomes "\frac{\varepsilon E^2}{4}" but in electromagnetic wave same amount is associated with magnetic field of induction so the average em energy per unit volume is "\frac{\varepsilon E^2}{2}."